OpenStudy (anonymous):

graph y=x-3. is the point (3,0) on the line?

OpenStudy (anonymous):

wheres the graph

OpenStudy (solomonzelman):

Your line is: \(\large\color{blue}{ y=x-3 }\). To find out if a point \(\large\color{blue}{ (3,0) }\) is on a line, you would simply plug in the point into the given line. For example \(\small\color{royalblue}{\rm (showing~how~to~plug~in~your~point) }\) if you wanted to find out, whether a point \(\large\color{blue}{(1,-2) }\) is on a line or not: \(\large\color{blue}{ y=x-3 }\) \(\large\color{blue}{ (-2)=(1)-3 }\) \(\large\color{blue}{-2=-2}\) and hence, a point \(\large\color{blue}{(1,-2) }\) is on the line.